OpenStudy (anonymous):

what is 3 1/4 + 4 1/3

OpenStudy (anonymous):

this is kinda the same as the last one but a lil more complex what is 3 in terms of 1/4?

OpenStudy (anonymous):

12/4, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so then 3+1/4=?

OpenStudy (anonymous):

i dont know im sorry

OpenStudy (anonymous):

ok how about this

OpenStudy (anonymous):

3 is 12/4. so what is \[\frac{ 12 }{ 4 }+\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

12/4

OpenStudy (anonymous):

13/4

OpenStudy (anonymous):

so i meant 13

OpenStudy (anonymous):

imagine if you had 12 quarters, right?

OpenStudy (anonymous):

12 quarters is 12/4 dollars, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then you add one more quarter, now you have 13/4 dollars

OpenStudy (anonymous):

so now that we know that 3 1/4=13/4, we're about half done with the question

OpenStudy (anonymous):

if 3 in terms of 1/4 is 12/4, can you figure out what 4 is in terms of 1/3?

OpenStudy (anonymous):

its 12/3.

OpenStudy (anonymous):

do you see how i did that, or do you want me to explain it?

OpenStudy (anonymous):

you had to see what they had in common right?

OpenStudy (anonymous):

i havent gotten there yet ive just been turning mixed fractions into improper

OpenStudy (anonymous):

do you know how 3=12/4?

OpenStudy (anonymous):

because since we have 3+1/4 in the first half of the question, we have to change that into just a fraction with no whole number, or else we wont be able to add the fractions

OpenStudy (anonymous):

so since its 3+1/4, i went with 4 for the denominator, so you have basically (3/1)=(x/4)

OpenStudy (anonymous):

do you know how to cross multiply?

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

ok so basically \[\frac{ 3 }{ 1}=\frac{ x }{ 4 }\]

OpenStudy (anonymous):

so multiply 1*x=3*4

OpenStudy (anonymous):

so x=12

OpenStudy (anonymous):

and plug that back in to x/4, so it equals 12/4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok so 3 1/4=13/4

OpenStudy (anonymous):

can you figure out what 4 1/3 is in terms of 1/3?

OpenStudy (anonymous):

using cross multiplication

OpenStudy (anonymous):

(4/1)=(x/3) solve for x

OpenStudy (anonymous):

4/3?

OpenStudy (anonymous):

well cross multiply it so you have 1*x=4*3

OpenStudy (anonymous):

then x=?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

now plug 12 back into the original equation x/3

OpenStudy (anonymous):

12/3

OpenStudy (anonymous):

do you see how i did that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

imagine if you have 4 apple pies

OpenStudy (anonymous):

and you cut each pie into 3 equal pieces

OpenStudy (anonymous):

how many pieces would you have?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

4 pies. 3 pieces.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

right. so 4 pieces in 3 pieces each is 12/3 pies

OpenStudy (anonymous):

because you have 12 pieces, and 3 pieces per pie

OpenStudy (anonymous):

lets try another example, lets say you have 5 dollars, and break it into quarters. how many quarters would you have?

OpenStudy (anonymous):

hey im sorry if this is annoying to you, if you want me to just answer the question you can tell me so, im just trying to help you understand it better.

OpenStudy (anonymous):

ryan played hockey for 3 3/4 hours on saturday and for 1 3/5 hours on sunday. how many hours did he pay hockey during these 2 days?

OpenStudy (anonymous):

my bed time is coming close an i have a 21 question test i have to get done. i think u for exampling it to me

OpenStudy (anonymous):

(3 1/4)+(4 1/3)=(13/4)+(13/3)=(39/12)+(52/12)=91/12

OpenStudy (anonymous):

thats for your first question

OpenStudy (anonymous):

(3 3/4)+(1 3/5)=(15/4)+(8/5)=(75/20)+(32/20)=107/20

OpenStudy (anonymous):

thats the second

OpenStudy (anonymous):

its not my answer choices

OpenStudy (anonymous):

for both of them?

OpenStudy (anonymous):

ohh they probably made them mixed fractions again

OpenStudy (anonymous):

on the 1st question my choices are 7 5/6, 7 1/2, 7 7/12, 7 4/7

OpenStudy (anonymous):

right so the first one is 7 1/2 and the second one is 5 7/20

OpenStudy (anonymous):

wait wait the first is 7 7/12

OpenStudy (anonymous):

my bad lol

OpenStudy (anonymous):

because 7=84/12, then 91/12= 7+7/12

OpenStudy (anonymous):

well, (7)+(7/12) to be more clear but im sure you get me

OpenStudy (anonymous):

the second one is 5 7/20 because 5=100/20 so 107/20=(5)+(7/20)

OpenStudy (anonymous):

good? :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

mrs jennings spent 14 1/6 hours at the gym last month. mr jennings spent 12 5/6 hours at the gym last month. how much more time did mrs jennings spend at the gym than mr jennings